# Trigonometry of partially truncated hyperbolic triangles and tetrahedra

### v2

@@ 2. Dimension two
Let $P$ be a pentagon with four right angles and a single ideal vertex, and let $\mathbf{x}$ be a positive light-like vector that determines a horosphere $S$ centered at the ideal vertex of $P$. Labeling the endpoints of the edge of $P$ opposite its ideal vertex as $\mathbf{w}_{0}$ and $\mathbf{w}_{1}$, for $i=0,1$ let $\gamma_{i}$ be the other edge of $P$ containing $\mathbf{w}_{i}$, and let $\mathbf{y}_{i}$ be a unit space-like vector in $\mathbb{R}^{3}$ orthogonal to the geodesic containing $\gamma_{i}$. Choose the $\mathbf{y}_{i}$ so that $\mathbf{y}_{i}\circ\mathbf{x}<0$ for each $i$. Equivalently, by Lemma 1.8, $\mathbf{y}_{i}$ is on the opposite side of $\mathbf{x}$ from the plane $\mathbf{u}\circ\mathbf{x}=0$ in $\mathbb{R}^{3}$. Since $\gamma_{i}$ and the ideal point of $P$ are on the same side of the geodesic containing $\gamma_{1-i}$ for each $i$, $\mathbf{y}_{0}\circ\mathbf{y}_{1}<0$ by [10].
Let us call $\mathbf{v}_{i}$ the endpoint of $\gamma_{i}$ not equal to $\mathbf{w}_{i}$, for $i=0,1$. An explicit formula for $\mathbf{v}_{i}$ is given by (4), with $\mathbf{y}$ there replaced by $\mathbf{y}_{i}$. As in the proofs of Theorem 3.2.7 and 3.2.8 of [10] we have the following explicit formula for $\mathbf{w}_{i}$:
$\mathbf{w}_{i}=\frac{-(\mathbf{y}_{0}\circ\mathbf{y}_{1})\mathbf{y}_{i}+\mathbf{y}_{i-1}}{\pm\sqrt{(\mathbf{y}_{0}\circ\mathbf{y}_{1})^{2}-1}},$
where “$+$” or “--” is chosen so that $\mathbf{w}_{i}$ is a positive vector. For, say, $i=0$ we thus have
@@
In passing from the first to the second equality we have fixed the sign choice “$+$” for the radical. This is the right choice since $\mathbf{y}_{0}\circ\mathbf{y}_{1}$ and the $\mathbf{x}\circ\mathbf{y}_{i}$ are all negative, and $\mathbf{w}_{0}\circ\mathbf{v}_{0}$ is as well.
@@ 3. Dimension three: transversals of truncated tetrahedra
$\mathbf{v}_{1}$$\mathbf{y}_{3}$$\mathbf{y}_{2}$$\mathbf{y}_{4}$$\mathbf{y}_{1}$$\mathbf{v}_{3}$
@@
The Hessian of each of $e^{s}e^{t}$ and $e^{-s}e^{-t}$ is a scalar multiple of $\left(\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right)$, and that of each of $e^{s}e^{-t}$ and $e^{-s}e^{t}$ is a multiple of $\left(\begin{smallmatrix}1&-1\\ -1&1\end{smallmatrix}\right)$. Each of these matrices has 00 and $2$ as eigenvalues and $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ and $\left(\begin{smallmatrix}1\\ -1\end{smallmatrix}\right)$ as eigenvectors, but the null eigenvector of either is associated to the eigenvalue $2$ of the other. It follows that $D(s,t)$, being a positive linear combination of the four functions, has positive-definite Hessian and therefore is strictly convex.
@@
Setting each of these equal to 00, subtracting bottom from top, and multiplying by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{+-}e^{2s_{0}}-C_{-+}e^{2t_{0}}$, giving $e^{2s_{0}}=e^{2t_{0}}C_{-+}/C_{+-}$ for the critical point $(s_{0},t_{0})$. Adding the two equations and multiplying through by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{++}e^{2s_{0}}e^{2t_{0}}-C_{--}$. After substituting for $e^{2s_{0}}$ we solve for $e^{2t_{0}}$, then plug the result back in to obtain the formulas of (6).
The sequence of inequalities in this result’s statement that starts with $0<C_{++}$ implies that $C_{--}/C_{++}$ is greater than both $C_{+-}/C_{-+}$ and its reciprocal $C_{-+}/C_{+-}$, which in turn implies that the quantities in parenthesis in (6) are each greater than $1$. Therefore if these inequalities hold then each of $s_{0}$ and $t_{0}$ is greater than 00. ∎
@@
Given the claim, Lemma 3.3 implies assertion (1) of this result. The fact that the coefficients $C_{\pm\pm}$ depend only on the $L_{ij}$ implies the same for the coordinates of the critical point $(s_{0},t_{0})$ of $D$, and hence for the absolute minimum value $D(s_{0},t_{0})$. Lemma 3.3 also gives that $(s_{0},t_{0})\in(0,\infty)^{2}$; because $\tilde{\lambda}_{ij}$ was parametrized pointing into $H_{i}$ with $\tilde{\lambda}_{ij}(0)\in P_{i}$, the fact that $s_{0}>0$ implies that the closest point $\tilde{\lambda}_{ij}(s_{0})$ to $\tilde{\lambda}_{kl}$ lies in $H_{i}$. Likewise, $\tilde{\lambda}_{kl}(t_{0})\in H_{k}$. But by swapping the roles of $i$ and $j$, and of $k$ and $l$—which only re-parametrizes $\tilde{\lambda}_{12}$ and/or $\tilde{\lambda}_{34}$—and re-running the argument above, we find that the closest point of $\tilde{\lambda}_{ij}$ to $\tilde{\lambda}_{kl}$ also lies in $H_{j}$. Therefore it lies in the edge $\lambda_{ij}$ of $\Delta$. Likewise, the closest point of $\tilde{\lambda}_{kl}$ to $\tilde{\lambda}_{ij}$ lies in the edge $\lambda_{kl}$. Therefore the geodesic arc joining these two points is the relevant transversal of $\Delta$, as asserted in (3) above.
@@
It remains to show that $T(x,y;a,b,c,d)$ is increasing in each of its last four variables. For this (again taking $i=1$, $j=2$, $k=3$, and $l=4$ after renumbering the $P_{i}$ if necessary), we claim that for any fixed $(s,t)$ the value $D(s,t)$—with itself depends on the values $x=L_{12}$, $y=L_{34}$, $a=L_{13}$, $b=L_{14}$, $c=L_{23}$, and $d=L_{24}$—increases with $d=L_{24}$ (taking the other $L_{ij}$ as fixed). To see this, we plug the values from (9) into the formula of (7). Dividing out by $\cosh s\cosh t$ yields:
@@
The value $d=L_{24}$ appears only once in this formula, in the numerator of the first summand above, with a positive sign. The claim is thus clear. It follows that the absolute minimum of $D(s,t)$, and hence also the transversal length $T(x,y;a,b,c,d)$, also increases with $d$. Since values of $T(x,y;a,b,c,d)$ are invariant under a transitive group action on $\{a,b,c,d\}$, the same then holds for $a$, $b$ and $c$.∎
@@ 4. Dimension three: partially truncated tetrahedra
For a fixed $k\in\{1,2,3\}$, suppose $P_{1},\ldots,P_{k}$ are pairwise disjoint and non-parallel planes, and $B_{k+1},\ldots,B_{4}$ are horoballs defined by positive light-like vectors $\mathbf{x}_{k+1},\ldots,\mathbf{x}_{4}$ such that the ideal points $[\mathbf{x}_{j}]$ are pairwise distinct and each not contained in the ideal boundary of any $P_{i}$. Further, suppose for each $i\leq k$ that there is a single half-space $H_{i}$ bounded by $P_{i}$ which contains $P_{j}$ for all $j\neq i\leq k$, and whose ideal boundary contains each ideal point $[\mathbf{x}_{j^{\prime}}]$.
For any fixed $i$, there is a unique plane $\widehat{P}_{i}$ that intersects $P_{j}$ at right angles for each $j\neq i\leq k$, and whose ideal boundary contains each ideal point $[\mathbf{x}_{j^{\prime}}]$ for $j^{\prime}\neq i>k$ (whence $\widehat{P}_{i}$ also intersects $B_{j^{\prime}}$ at a right angle).
@@
    For $i>k$, if the ideal boundary of $\widehat{P}_{i}$ does not contain $[\mathbf{x}_{i}]$ then there is a single half-space $\widehat{H}_{i}$ bounded by $\widehat{P}_{i}$ such that for each $j\leq k$, $\widehat{H}_{i}$ contains the entire geodesic perpendicular to $P_{j}$ in the direction of $\mathbf{x}_{i}$, and for each $j^{\prime}\neq i>k$, the entire geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j^{\prime}}$.
@@
As in the proof of Lemma 3.1, $\widehat{P}_{i}$ contains the mutual perpendicular geodesic to $P_{j}$ and $P_{j^{\prime}}$, for any distinct $j$ and $j^{\prime}\leq k$ and not $i$, so it meets each such $P_{j}$ at right angles. It contains the ideal points $[\mathbf{x}_{j}]$, $j\neq i>k$, by construction. Moreover, for $j,j^{\prime}\neq i$, if $j\leq k$ and $j^{\prime}>k$ then $\widehat{P}_{i}$ contains the geodesic perpendicular to $P_{j}$ in the direction of $\mathbf{x}_{j^{\prime}}$ described in Lemma 1.8; and if $j,j^{\prime}>k$ then $\widehat{P}_{i}$ contains the geodesic between $\mathbf{x}_{j}$ and $\mathbf{x}_{j^{\prime}}$ described in Lemma 1.4. As Remarked below Lemma 1.3, this implies that $\widehat{P}_{i}$ intersects each $B_{j}$ perpendicularly.
@@
Now take $i>k$ and suppose that the ideal boundary of $\widehat{P}_{i}$ does not contain $[\mathbf{x}_{i}]$. For any $j\leq k$, the geodesic perpendicular to $P_{j}$ and in the direction of $\mathbf{x}_{i}$ is described as $\gamma_{j}(t)=e^{-h}\cosh t\,\mathbf{x}_{i}+e^{-t}\mathbf{y}_{j}$ by Lemma 1.8, where $e^{h}=-\mathbf{x}_{i}\circ\mathbf{y}_{j}$. For a unit normal $\mathbf{z}_{i}$ to $\widehat{P}_{i}$ and any $t\in\mathbb{R}$, $\gamma_{j}(t)\circ\mathbf{z}_{i}=e^{-h}\cosh t\,\mathbf{x}_{i}\circ\mathbf{z}_{i}$ has the same sign as $\mathbf{x}_{i}\circ\mathbf{z}_{i}$. Choosing $\mathbf{z}_{i}$ to be the outward unit normal to the half-space $\widehat{H}_{i}$ containing $\mathbf{x}_{i}$ in its ideal boundary, it follows that $\gamma_{j}(t)$ is contained in $\widehat{H}_{i}$ for all $t\in\mathbb{R}$. For any $j^{\prime}\neq i>k$, the geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j^{\prime}}$ has the form $\gamma_{j^{\prime}}(t)=\frac{1}{2}e^{-d/2}\left(e^{t}\,\mathbf{x}_{i}+e^{-t}\,\mathbf{x}_{j^{\prime}}\right)$ by Lemma 1.4, where $e^{d}=-\frac{1}{2}\mathbf{x}_{i}\circ\mathbf{x}_{j^{\prime}}$. Thus for $\mathbf{z}_{i}$ as above, since $\mathbf{x}_{j^{\prime}}$ is in the ideal boundary of $\widehat{P}_{i}$, $\gamma_{j^{\prime}}(t)\circ\mathbf{z}_{i}=\frac{1}{2}e^{-d/2}e^{t}\,\mathbf{x}_{i}\circ\mathbf{z}_{i}<0$. Therefore again $\gamma_{j^{\prime}}(t)\in\widehat{H}_{i}$ for all $t$. This establishes the Lemma’s conclusion (2). ∎
@@
For a fixed $k\in\{1,2,3\}$, suppose $P_{1},\ldots,P_{k}$ are pairwise disjoint and non-parallel planes, and $B_{k+1},\ldots,B_{4}$ are horoballs defined by positive light-like vectors $\mathbf{x}_{k+1},\ldots,\mathbf{x}_{4}$ such that the ideal points $[\mathbf{x}_{j}]$ are pairwise distinct and each not contained in the ideal boundary of any $P_{i}$. Further, suppose for each $i\leq k$ that there is a single half-space $H_{i}$ bounded by $P_{i}$ which contains $P_{j}$ for all $j\neq i\leq k$, and whose ideal boundary contains each ideal point $[\mathbf{x}_{j^{\prime}}]$.
@@
In all cases, for each $i>k$, we say the projective class $[\mathbf{x}_{i}]$ of $\mathbf{x}_{i}$ is an ideal vertex of $\widehat{P}_{i}$. For each $i<j\leq k$, the internal edge $\lambda_{ij}$ of $\Delta$ is the shortest arc in $\mathbb{H}^{3}$ joining $P_{i}$ to $P_{j}$; for $i\leq k<j$, the ray edge $\lambda_{ij}$ is the geodesic ray perpendicular to $P_{i}$ in the direction of $\mathbf{x}_{j}$; and for $k<i<j$, the bi-infinite edge $\lambda_{ij}$ is the geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j}$. The edge opposite $\lambda_{ij}$ is $\lambda_{i^{\prime}j^{\prime}}$, where $i^{\prime}<j^{\prime}\in\{1,2,3,4\}-\{i,j\}$. For each $i$, the internal face opposite $P_{i}$ or $\mathbf{x}_{i}$ is the polygon in $\Delta\cap\widehat{P}_{i}$ bounded by the internal edges $\lambda_{jj^{\prime}}$, for each pair $j<j^{\prime}\in\{1,2,3,4\}-\{i\}$, and arcs of the $P_{j}$, $j\neq i\leq k$. The non-internal faces and edges of $\Delta$ are external. Each of these is entirely contained in $P_{i}$ for some $i\leq k$.

### v3

@@ 2. Dimension two
Let $P$ be a pentagon with four right angles and a single ideal vertex, and let $\mathbf{x}$ be a positive light-like vector that determines a horosphere $S$ centered at the ideal vertex of $P$. Labeling the endpoints of the edge $\delta$ of $P$ opposite its ideal vertex as $\mathbf{w}_{0}$ and $\mathbf{w}_{1}$, for $i=0,1$ let $\gamma_{i}$ be the other edge of $P$ containing $\mathbf{w}_{i}$, and let $\mathbf{y}_{i}$ be a unit space-like vector in $\mathbb{R}^{3}$ orthogonal to the geodesic containing $\gamma_{i}$. Choose the $\mathbf{y}_{i}$ so that $\mathbf{y}_{i}\circ\mathbf{x}<0$ for each $i$. Equivalently, by Lemma 1.8, $\mathbf{y}_{i}$ is the outward normal to the half-space $H_{i}$ bounded by $\gamma_{i}$ and containing $\mathbf{x}$ in its ideal boundary. This half-space also contains $\gamma_{1-i}$ in its interior; therefore by construction, $\mathbf{y}_{0}$ and $\mathbf{y}_{1}$ are oppositely-pointing tangent vectors to $\delta$, so by Lemma 1.9(2), $\mathbf{y}_{0}\circ\mathbf{y}_{1}<0$.
Lemma 1.10 gives the formula below for the $\mathbf{w}_{i}$, bearing in mind that for each $i$, $\mathbf{w}_{i}\in H_{1-i}$.
$\mathbf{w}_{i}=\frac{-(\mathbf{y}_{0}\circ\mathbf{y}_{1})\mathbf{y}_{i}+\mathbf{y}_{1-i}}{\sqrt{(\mathbf{y}_{0}\circ\mathbf{y}_{1})^{2}-1}},$
Let us call $\mathbf{v}_{i}$ the endpoint of $\gamma_{i}$ not equal to $\mathbf{w}_{i}$, for $i=0,1$. An explicit formula for $\mathbf{v}_{i}$ is given by (5), with $\mathbf{y}$ there replaced by $\mathbf{y}_{i}$. For, say, $i=0$ we thus have
@@
Proposition 2.3.
Let $\Delta$ be an ideal triangle in $\mathbb{H}^{2}$ and $B_{1},B_{2},B_{3}$ be horoballs, one centered at each ideal vertex of $\Delta$. For $i\in\{1,2,3\}$, let $\theta_{i}$ be the length of the horocyclic arc $S_{i}\cap\Delta$, where $S_{i}=\partial B_{i}$, and for each $i<j$ let $d_{ij}$ be the signed distance between $B_{i}$ and $B_{j}$ along the geodesic joining their centers (with a negative sign if the horoballs overlap). Then:
  • (Law of Sines) ${\frac{\theta_{1}}{e^{d_{23}}}=\frac{\theta_{2}}{e^{d_{13}}}=\frac{\theta_{3}}{e^{d_{12}}}}$
  • (First Law of Cosines) ${\theta_{1}=\sqrt{\frac{e^{d_{23}}}{e^{d_{12}}e^{d_{13}}}}}$
  • (Second Law of Cosines) ${e^{d_{23}}=\frac{1}{\theta_{2}\theta_{3}}}$ [This was proved as Lemma 3.3 of [9].]
Proof.
Let $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, and $\mathbf{x}_{3}$ be positive light-like vectors respectively determining $B_{1}$, $B_{2}$, and $B_{3}$, and for each $i<j$ let $\lambda_{ij}(t)=\frac{1}{2}e^{-d_{ij}/2}\left(e^{t}\mathbf{x}_{i}+e^{-t}\mathbf{x}_{j}\right)$ be the geodesic from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ as in Lemma 1.4. By that result, the points of intersection $\mathbf{u}_{12}=\lambda_{12}\cap S_{1}$ and $\mathbf{u}_{13}=\lambda_{13}\cap S_{1}$ are given as:
$\mathbf{u}_{12}=\frac{1}{2}\left(\mathbf{x}_{1}+e^{-d_{12}}\mathbf{x}_{2}\right)\quad\mbox{and}\quad\mathbf{u}_{13}=\frac{1}{2}\left(\mathbf{x}_{1}+e^{-d_{13}}\mathbf{x}_{3}\right).$
Using the fact that for $i\neq j$, $\mathbf{x}_{i}\circ\mathbf{x}_{j}=-2e^{d_{ij}}$ (again by Lemma 1.4), we obtain
$\mathbf{u}_{12}\circ\mathbf{u}_{13}=-1-\frac{e^{d_{23}}}{2e^{d_{12}}e^{d_{13}}}\quad\Rightarrow\quad\theta_{1}=\sqrt{\frac{e^{d_{23}}}{e^{d_{12}}e^{d_{13}}}}.$
The formula for $\theta_{1}$ above comes from (2). It is the “First Law of Cosines” above. Formulas for $\theta_{2}$ and $\theta_{3}$ are completely analogous, and from these we obtain the “Law of Sines”:
$\frac{\theta_{1}}{e^{d_{23}}}=\sqrt{\frac{1}{e^{d_{12}}e^{d_{13}}e^{d_{23}}}}=\frac{\theta_{2}}{e^{d_{13}}}=\frac{\theta_{3}}{e^{d_{12}}}$
For the “Second Law of Cosines” above we simply multiply the formulas for $\theta_{2}$ and $\theta_{3}$ and solve for $e^{d_{23}}$. ∎
@@ 3. Dimension three: transversals of truncated tetrahedra
$\mathbf{y}_{1}$$\mathbf{y}_{4}$$\mathbf{y}_{2}$$\mathbf{y}_{3}$$\mathbf{v}_{1}$$\mathbf{v}_{3}$
@@
The Hessian of each of $e^{s}e^{t}$ and $e^{-s}e^{-t}$ is a scalar multiple of $\left(\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right)$, and that of each of $e^{s}e^{-t}$ and $e^{-s}e^{t}$ is a multiple of $\left(\begin{smallmatrix}1&-1\\ -1&1\end{smallmatrix}\right)$. Each of these matrices has 0 and $2$ as eigenvalues and $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ and $\left(\begin{smallmatrix}1\\ -1\end{smallmatrix}\right)$ as eigenvectors, but the null eigenvector of either is associated to the eigenvalue $2$ of the other. It follows that $D(s,t)$, being a positive linear combination of the four functions, has positive-definite Hessian and therefore is strictly convex.
@@
Setting each of these equal to 0, subtracting bottom from top, and multiplying by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{+-}e^{2s_{0}}-C_{-+}e^{2t_{0}}$, giving $e^{2s_{0}}=e^{2t_{0}}C_{-+}/C_{+-}$ for the critical point $(s_{0},t_{0})$. Adding the two equations and multiplying through by $e^{s_{0}}e^{t_{0}}/2$ yields $0=C_{++}e^{2s_{0}}e^{2t_{0}}-C_{--}$. After substituting for $e^{2s_{0}}$ we solve for $e^{2t_{0}}$, then plug the result back in to obtain the formulas of (7).
The sequence of inequalities in this result’s statement that starts with $0<C_{++}$ implies that $C_{--}/C_{++}$ is greater than both $C_{+-}/C_{-+}$ and its reciprocal $C_{-+}/C_{+-}$, which in turn implies that the quantities in parenthesis in (7) are each greater than $1$. Therefore if these inequalities hold then each of $s_{0}$ and $t_{0}$ is greater than 0. ∎
@@
Given the claim, Lemma 3.3 implies assertion (1) of this result. The fact that the coefficients $C_{\pm\pm}$ depend only on the $L_{ij}$ implies the same for the coordinates of the critical point $(s_{0},t_{0})$ of $D$, and hence for the absolute minimum value $D(s_{0},t_{0})$. Lemma 3.3 also gives that $(s_{0},t_{0})\in(0,\infty)^{2}$; because $\tilde{\lambda}_{ij}$ was parametrized pointing into $H_{i}$ with $\tilde{\lambda}_{ij}(0)\in P_{i}$, the fact that $s_{0}>0$ implies that the closest point $\tilde{\lambda}_{ij}(s_{0})$ to $\tilde{\lambda}_{kl}$ lies in $H_{i}$. Likewise, $\tilde{\lambda}_{kl}(t_{0})\in H_{k}$. But by swapping the roles of $i$ and $j$, and of $k$ and ll—which only re-parametrizes $\tilde{\lambda}_{12}$ and/or $\tilde{\lambda}_{34}$—and re-running the argument above, we find that the closest point of $\tilde{\lambda}_{ij}$ to $\tilde{\lambda}_{kl}$ also lies in $H_{j}$. Therefore it lies in the edge $\lambda_{ij}$ of $\Delta$. Likewise, the closest point of $\tilde{\lambda}_{kl}$ to $\tilde{\lambda}_{ij}$ lies in the edge $\lambda_{kl}$. Therefore the geodesic arc joining these two points is the relevant transversal of $\Delta$, as asserted in (3) above.
@@
It remains to show that $T(x,y;a,b,c,d)$ is increasing in each of its last four variables. For this (again taking $i=1$, $j=2$, $k=3$, and $l=4$ after renumbering the $P_{i}$ if necessary), we claim that for any fixed $(s,t)$ the value D(s,t)D(s,t)—with itself depends on the values $x=L_{12}$, $y=L_{34}$, $a=L_{13}$, $b=L_{14}$, $c=L_{23}$, and $d=L_{24}$—increases with $d=L_{24}$ (taking the other $L_{ij}$ as fixed). To see this, we plug the values from (10) into the formula of (8). Dividing out by $\cosh s\cosh t$ yields:
@@
The value $d=L_{24}$ appears only once in this formula, in the numerator of the first summand above, with a positive sign. The claim is thus clear. It follows that the absolute minimum of $D(s,t)$, and hence also the transversal length $T(x,y;a,b,c,d)$, also increases with $d$. Since values of $T(x,y;a,b,c,d)$ are invariant under a transitive group action on $\{a,b,c,d\}$, the same then holds for $a$, $b$ and cc.∎
@@ 4. Dimension three: partially truncated tetrahedra
For a fixed $k\in\{1,2,3\}$, suppose $P_{1},\ldots,P_{k}$ are pairwise disjoint and non-parallel planes, and $B_{k+1},\ldots,B_{4}$ are horoballs defined by positive light-like vectors $\mathbf{x}_{k+1},\ldots,\mathbf{x}_{4}$ such that the ideal points $\mathbf{x}_{j}]$ are pairwise distinct and each not contained in the ideal boundary of any $P_{i}$. Further, suppose for each $i\leq k$ that there is a single half-space $H_{i}$ bounded by $P_{i}$ which contains $P_{j}$ for all $j\neq i\leq k$, and whose ideal boundary contains each ideal point $\mathbf{x}_{j^{\prime}}]$.
For any fixed $i$, there is a unique plane $\widehat{P}_{i}$ that intersects $P_{j}$ at right angles for each $j\neq i\leq k$, and whose ideal boundary contains each ideal point $\mathbf{x}_{j^{\prime}}]$ for $j^{\prime}\neq i>k$ (whence $\widehat{P}_{i}$ also intersects $B_{j^{\prime}}$ at a right angle).
@@
    For $i>k$, if the ideal boundary of $\widehat{P}_{i}$ does not contain $\mathbf{x}_{i}]$ then there is a single half-space $\widehat{H}_{i}$ bounded by $\widehat{P}_{i}$ such that for each $j\leq k$, $\widehat{H}_{i}$ contains the entire geodesic perpendicular to $P_{j}$ in the direction of $\mathbf{x}_{i}$, and for each $j^{\prime}\neq i>k$, the entire geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j^{\prime}}$.
@@
As in the proof of Lemma 3.1, $\widehat{P}_{i}$ contains the mutual perpendicular geodesic to $P_{j}$ and $P_{j^{\prime}}$, for any distinct $j$ and $j^{\prime}\leq k$ and not $i$, so it meets each such $P_{j}$ at right angles. It contains the ideal points $\mathbf{x}_{j}]$, $j\neq i>k$, by construction. Moreover, for $j,j^{\prime}\neq i$, if $j\leq k$ and $j^{\prime}>k$ then $\widehat{P}_{i}$ contains the geodesic perpendicular to $P_{j}$ in the direction of $\mathbf{x}_{j^{\prime}}$ described in Lemma 1.8; and if $j,j^{\prime}>k$ then $\widehat{P}_{i}$ contains the geodesic between $\mathbf{x}_{j}$ and $\mathbf{x}_{j^{\prime}}$ described in Lemma 1.4. As Remarked below Lemma 1.3, this implies that $\widehat{P}_{i}$ intersects each $B_{j}$ perpendicularly.
@@
Now take $i>k$ and suppose that the ideal boundary of $\widehat{P}_{i}$ does not contain $\mathbf{x}_{i}]$. For any $j\leq k$, the geodesic perpendicular to $P_{j}$ and in the direction of $\mathbf{x}_{i}$ is described as $\gamma_{j}(t)=e^{-h}\cosh t\,\mathbf{x}_{i}+e^{-t}\mathbf{y}_{j}$ by Lemma 1.8, where $e^{h}=-\mathbf{x}_{i}\circ\mathbf{y}_{j}$. For a unit normal $\mathbf{z}_{i}$ to $\widehat{P}_{i}$ and any $t\in\mathbb{R}$, $\gamma_{j}(t)\circ\mathbf{z}_{i}=e^{-h}\cosh t\,\mathbf{x}_{i}\circ\mathbf{z}_{i}$ has the same sign as $\mathbf{x}_{i}\circ\mathbf{z}_{i}$. Choosing $\mathbf{z}_{i}$ to be the outward unit normal to the half-space $\widehat{H}_{i}$ containing $\mathbf{x}_{i}$ in its ideal boundary, it follows that $\gamma_{j}(t)$ is contained in $\widehat{H}_{i}$ for all $t\in\mathbb{R}$. For any $j^{\prime}\neq i>k$, the geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j^{\prime}}$ has the form $\gamma_{j^{\prime}}(t)=\frac{1}{2}e^{-d/2}\left(e^{t}\,\mathbf{x}_{i}+e^{-t}\,\mathbf{x}_{j^{\prime}}\right)$ by Lemma 1.4, where $e^{d}=-\frac{1}{2}\mathbf{x}_{i}\circ\mathbf{x}_{j^{\prime}}$. Thus for $\mathbf{z}_{i}$ as above, since $\mathbf{x}_{j^{\prime}}$ is in the ideal boundary of $\widehat{P}_{i}$, $\gamma_{j^{\prime}}(t)\circ\mathbf{z}_{i}=\frac{1}{2}e^{-d/2}e^{t}\,\mathbf{x}_{i}\circ\mathbf{z}_{i}<0$. Therefore again $\gamma_{j^{\prime}}(t)\in\widehat{H}_{i}$ for all $t$. This establishes the Lemma’s conclusion (2). ∎
@@
For a fixed $k\in\{1,2,3\}$, suppose $P_{1},\ldots,P_{k}$ are pairwise disjoint and non-parallel planes, and $B_{k+1},\ldots,B_{4}$ are horoballs defined by positive light-like vectors $\mathbf{x}_{k+1},\ldots,\mathbf{x}_{4}$ such that the ideal points $\mathbf{x}_{j}]$ are pairwise distinct and each not contained in the ideal boundary of any $P_{i}$. Further, suppose for each $i\leq k$ that there is a single half-space $H_{i}$ bounded by $P_{i}$ which contains $P_{j}$ for all $j\neq i\leq k$, and whose ideal boundary contains each ideal point $\mathbf{x}_{j^{\prime}}]$.
@@
In all cases, for each $i>k$, we say the projective class $\mathbf{x}_{i}]$ of $\mathbf{x}_{i}$ is an ideal vertex of $\widehat{P}_{i}$. For each $i<j\leq k$, the internal edge $\lambda_{ij}$ of $\Delta$ is the shortest arc in $\mathbb{H}^{3}$ joining $P_{i}$ to $P_{j}$; for $i\leq k<j$, the ray edge $\lambda_{ij}$ is the geodesic ray perpendicular to $P_{i}$ in the direction of $\mathbf{x}_{j}$; and for $k<i<j$, the bi-infinite edge $\lambda_{ij}$ is the geodesic joining $\mathbf{x}_{i}$ to $\mathbf{x}_{j}$. The edge opposite $\lambda_{ij}$ is $\lambda_{i^{\prime}j^{\prime}}$, where $i^{\prime}<j^{\prime}\in\{1,2,3,4\}-\{i,j\}$. For each $i$, the internal face opposite $P_{i}$ or $\mathbf{x}_{i}$ is the polygon in $\Delta\cap\widehat{P}_{i}$ bounded by the internal edges $\lambda_{jj^{\prime}}$, for each pair $j<j^{\prime}\in\{1,2,3,4\}-\{i\}$, and arcs of the $P_{j}$, $j\neq i\leq k$. The non-internal faces and edges of $\Delta$ are external. Each of these is entirely contained in $P_{i}$ for some $i\leq k$.